OpenStudy (lilmike234):

What is the value of a fourth term in a geometric sequence for which a1=10 and r=.5

OpenStudy (solomonzelman):

\(\LARGE a_n=a_1\times {\rm r}^{\rm n -1}\)

OpenStudy (solomonzelman):

this is for any geometric sequence.

OpenStudy (solomonzelman):

Now, for the fourth term of any geometric sequence it will read: \(\LARGE a_4=a_1\times {\rm r}^{\rm 4 -1}\) \(\LARGE a_4=a_1\times {\rm r}^{3}\)

OpenStudy (solomonzelman):

knowing your a1=10, and r=5, plug that into: \(\LARGE a_4=a_1\times {\rm r}^{3}\) and then calculate the \(a_4\).

OpenStudy (lilmike234):

So the answer would be 1250?

OpenStudy (solomonzelman):

lets see \(a_4=10\times 5^3\) \(a_4=10\times 125\) \(a_4=1250\) yes, your answer is correct:)

OpenStudy (lilmike234):

The answer was 1.25...